If the numerator and denominator of a fraction are both decreased by one the fraction becomes 2/3. If the numerator and denominator are both increased by one the fraction becomes 3/4. Find the original fraction
Answer by stanbon(75887)

If the numerator and denominator of a fraction are both decreased by one the fraction becomes 2/3. If the numerator and denominator are both increased by one the fraction becomes 3/4. Find the original fraction
-----------------------------------------------
Let the fraction be x/y
---
Equations:
(x-1)/(y-1) = 2/3
(x+1)/(y+1) = 3/4
--------
Rearrange:
3x-3=2y-2
4x+4=3y+3
------------
Rearrange:
3x - 2y = 1
4x - 3y = -1
----------------------
Multiply 1st by 3
Multiply 2nd by 2
--------
9x - 6y = 2
8x - 6y = -2
-------------------------
Subtract and solve for "x":
x = 4
---
solve for "y":
3x-2y = 1
12-2y = 1
2y = 11
y = 11/2
---
Solution: (11/2,4)
========================
Cheers,
Stan H.